OpenStudy (anonymous):

lim x-> pi/4 of (1-tanx)/(sinx-cosx)

terenzreignz (terenzreignz):

Did you try multiplying by conjugates?

OpenStudy (anonymous):

No I multiplied by cosx?

terenzreignz (terenzreignz):

Why, though?

OpenStudy (anonymous):

Because tanx becomes sinx/cosx so to cancel the denominator I multiplied by cos x so it becomes (cosx-sinx)/cosx•sinx-cosx and then the cosines equal 1 and as well do the sines so its -1/-cosx

OpenStudy (anonymous):

But the answer is supposed to be -square root of 2 but I got positive and don't know what i did wrong

terenzreignz (terenzreignz):

Fair enough, hang on.

terenzreignz (terenzreignz):

\[\Large \lim_{x\rightarrow \frac \pi 4}\frac{1-\tan(x)}{\sin(x) - \cos(x)}\]

terenzreignz (terenzreignz):

You missed a tiny detail, it would seem ^_^ \[\Large \frac{\cos(x)}{\cos(x)}\cdot \frac{1-\tan(x)}{\sin(x) - \cos(x)}= \frac{\cos(x)-\sin(x)}{\cos(x)[\sin(x) - \cos(x)]}\]

terenzreignz (terenzreignz):

See it now? \[\LARGE = \frac{\cos(x) - \sin(x)}{\cos(x)[\cos(x) -\sin(x)]\color{red}{(-1)}}\]

OpenStudy (anonymous):

Welcome to OpenStudy! Message me if you have any questions!

terenzreignz (terenzreignz):

Because these two are negatives of each other: \[\LARGE\frac{\color{red}{\cos(x)-\sin(x)}}{\cos(x)[\color{blue}{\sin(x) - \cos(x)}]}\]

OpenStudy (anonymous):

I'm new to this website but for some reason I can't see all of what you wrote

OpenStudy (anonymous):

Where you put you missed a tiny detail I can't see the whole problem and where you put see it now as well

terenzreignz (terenzreignz):

@Daisytoolovely sorry, was distracted... basically cos(x) - sin(x) = -[sin(x) - cos(x)] So that in effect, cos(x) - sin(x) / sin(x) - cos(x) = -1